What is the probability that the first child of individuals III-1 and III-2 will have the disorder?

The answer is 1/18 but i am not sure how you get to that answer.
Answer #1

Plas). E

The I-1 and I-2 are unaffected but have an affected son (II-1). So, both of them must be carriers (genotype-Dd). II-2 is D_, with a 1/3 chance of being DD and 2/3 chance of being Dd. II-5 and II-6 are both Dd because they are unaffected but have an affected daughter (III-3). The III-1 is D_, DD=1/2 and Dd=1/2. III-2 is D_, DD=1/3 and Dd=2/3. The IV-1 has chance of being dd is ¼.

So, the probability of the disorder of first child of individuals III-1 and III-2 is (2/3 X 1/2) X 2/3 X 1/4 = 1/18